OpenStudy (anonymous):

A circle has an area of 113.04 units2 and a circumference of 37.68 units. If the radius is 6 units, what can be said about the relationship between the area and the circumference? (Use 3.14 for pi.) A.The ratio of the area to the circumference is equal to half the radius. B.The ratio of the area to the circumference is equal to twice the radius. C.The ratio of the area to the circumference is equal to the square root of the radius. E.The ratio of the area to the circumference is equal to the radius squared.

OpenStudy (campbell_st):

well you need to recognise the ratio of Area to Circumference is \[\frac{Area}{Circumference} = \frac{113.04}{37.68}\] simplify the fraction for the answer the more abstract approach is using formula \[\frac{Area}{Circumference} = \frac{\pi r^2}{2\pi r}\] again, simplify the fraction

OpenStudy (anonymous):

i know what they both are

OpenStudy (anonymous):

sorry, but you didnt need to do that. I just didnt understand the question. It was kind of confusing me. Circumference = radius time 2 times pi. Area = radius squared times pi.

OpenStudy (anonymous):

@ParthKohli @texaschic101 @paki @undeadknight26

OpenStudy (anonymous):

Can any of you guy's help me???? @ParthKohli @texaschic101 @paki @undeadknight26

OpenStudy (campbell_st):

its about simplifying the fractions... that was mentioned in the post formulae... simplify by removing common factors \[\frac{\pi r^2}{2\pi r} = ?\] now confirm it with the measurements \[\frac{113.04}{37.68} = ?\] what conclusion can you make, then does it match you answer choices